OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)? y = 2x – 8 y = – 2x + 9 y = 2x – 11 y = – 2x + 6

OpenStudy (anonymous):

well first find the slope

OpenStudy (anonymous):

that will knock two answers off your list

OpenStudy (anonymous):

y - (-7) = (-1-(-7))/(5-2) (x-2) y+7 = 6/3 (x-2) y+7 = 2(x-2) y = 2x -4-7 y = 2x _11

OpenStudy (anonymous):

He's doing it all together but basically hes doing this \[m _{L}=(y _{2}-y _{1})/(x _{2}-x _{1})\]

OpenStudy (anonymous):

then usng \[y-y _{1}=m _{L}(x-x _{1})\]